Could someone please help with this problem? It is a question on my homework assignment that doesn't have any examples in our book and I have no idea how to do it.
The ratio of a side length of triangle A to a corresponding side length of triangle B is 5:8. Triangle A has a side length of 18 centimeters. Find the corresponding side length of triangle B.

Answer by scott8148(6628)

A/B = 5/8 = 18/x